OpenStudy (tmagloire1):

Find the x-coordinates of any relative extrema and inflection point(s) for the function f(x) = 3x^(1/3) + 6x^(4/3). You must justify your answer using an analysis of f ′(x) and f ′′(x).

OpenStudy (tmagloire1):

@phi @Directrix

OpenStudy (tmagloire1):

@amistre64

OpenStudy (amistre64):

define f' and f''

OpenStudy (tmagloire1):

f' = (1+8 x)/x^(2/3) f'' = (2 (-1+4 x))/(3 x^(5/3))

OpenStudy (amistre64):

f = 3x^(1/3) + 6x^(4/3) f' = 3/3 x^(-2/3) + 6*4/3 x^(1/3) f'' = 3/3* -2/3 x^(-5/3) + 6*4/3*1/3 x^(-2/3) how do we analyse them? what defines the critical points?

OpenStudy (tmagloire1):

Critical points are when you set them = 0 and see where they change direction?

OpenStudy (amistre64):

0 is only one of the steps ... also seeing where they are undefined is another

OpenStudy (tmagloire1):

How do you find where they're undefined?

OpenStudy (amistre64):

what makes a fraction undefined?

OpenStudy (amistre64):

x^(-n) = 1/x^n but this is not defined for what value of x?

OpenStudy (tmagloire1):

anything divided by 0

OpenStudy (amistre64):

then in this case that is going to be our undefineds

OpenStudy (tmagloire1):

How does that relate to the problem though?

OpenStudy (amistre64):

that is how you "Find the x-coordinates of any relative extrema and inflection point(s) for the function f(x)" f' and f'' = 0 or undefined ...

OpenStudy (tmagloire1):

Oh ok so is there a different way to find extrema and inflection points?

OpenStudy (amistre64):

precisely .. no graphing utility maybe, but by eye .. no

OpenStudy (tmagloire1):

so then extrema and inflection points are the same thing?

OpenStudy (amistre64):

no extrema are your highest/lowest ponts inflection is where slope changes direction

OpenStudy (amistre64):

 extrema, A and B inflection, M